- Press Release
- Oct 3, 2022
NASA Hubble Space Telescope Daily Report #4916
HUBBLE SPACE TELESCOPE DAILY REPORT #4916
PERIOD COVERED: 5am August 24 – 5am August 25, 2009 (DOY 236/09:00z-237/09:00z)
ACS CCD Monitoring and Calibration for WFC3
This program is a smaller version of our routine CCD monitoring program, designed to run throughout SMOV, after which our regular Cycle 17 CAL proposal will begin. This program obtains the bias and dark frames needed to generate reference files for calibrating science data, and allows us to monitor detector noise and the growth of hot pixels.
COS FUV Optics Alignment and Focus
After FUV detector functionality has been confirmed in COS25 (program 11483) and the initial focus updates determined in COS08 and COS 09 (programs 11468 and 11469), a sequence of 13 FUV focus-sweep exposures of a sharp-lined external target will be made with each grating (4 orbits per grating) to perform a fine-focus sweep. After the data are analyzed, a patchable constant SMS update of OSM1 focus for each grating will be uplinked. A verification visit will be executed after the uplink. Observations require high S/N and should be taken in TIME-TAG (FLASH=YES) mode.
COS FUV Initial On-Orbit Turn-On
This proposal specifies the procedure for SMOV initial HV turn-on and ramp-up of the COS FUV detector. (The FUV will have been commanded to its Operate state to support execution of proposal 11353.) The procedure is detailed in the Observing Description, but in summary, the following is done: The initial transition from FUV Operate to HVLow is broken into two parts, with a gap of 4 hours between turning on the HV and ramping to the HVLow (SAA) voltage. This will be followed by 5 cycles of HV ramp-up and return to HVLow. Cycles will ramp up to successively higher (magnitude) voltage, with the fifth cycle going to the nominal operating values. There will be a gap of at least 4 hours between cycles. All HV ramp-up will be done at 10 sec per HV “step”. The step rate and cycle voltage values (for Segments A and B) must be patched in FSW in each cycle prior to the HV ramp commanding. Memory monitors will be set on the patched memory locations. Immediately after any HV commanding, and 4 hours after ramp-up commanding, the DCE memory will be dumped. Immediately after HV ramp-up commanding higher than HVLow, short DARK & WAVE exposures will be obtained. Visits 01 and 02, and all the subsequent even numbered visits (the ones 4 hours after HV ramp-ups), end with NSSC-1 COS event flag 3 being set. If the flag remains set, subsequent FUV commanding will be skipped. Thus, Operations Requests must be in place to clear the flag prior to those subsequent visits. Real-time monitoring of the telemetry will be used to guide the decisions whether or not to clear the flag. The final visit (13) provides a scheduled final opportunity to clear flag 3, and if the flag is cleared, initiates nominal FUV HV commanding and requests a DARK exposure.
CCD Dark Monitor Part 1
Monitor the darks for the STIS CCD.
CCD Bias Monitor-Part 1
Monitor the bias in the 1×1, 1×2, 2×1, and 2×2 bin settings at gain=1, and 1×1 at gain = 4, to build up high-S/N superbiases and track the evolution of hot columns.
STIS CCD Spectroscopic Flats C17
Obtain pixel-to-pixel lamp flat fields for the STIS CCD in spectroscopic mode.
STIS-20 NUV MAMA Dark Monitor
The STIS NUV-MAMA dark current is dominated by a phosphorescent glow from the detector window. Meta-stable states in this window are populated by cosmic ray impacts, which, days later, can be thermally excited to an unstable state from which they decay, emitting a UV photon. The equilibrium population of these meta-stable states is larger at lower temperatures; so warming up the detector from its cold safing will lead to a large, but temporary, increase in the dark current.
To monitor the decay of this glow, and to determine the equilibrium dark current for Cycle 17, four 1380s NUV-MAMA ACCUM mode darks should be taken each week during the SMOV period. Once the observed dark current has reached an approximate equilibrium with the mean detector temperature, the frequency of this monitor can be reduced to one pair of darks per week.
STIS-26 MAMA Image Stability
The maximum thermal motion of the MAMA detectors occurs in the first portion of the orbit immediately following a large angle maneuver leading to maximum external changes on the portion of axial bay closest to the STIS instrument. By the second orbit on the same target, the thermal motions settle down to a significant displacement right after target rise, a possible change later in the orbit due to sun/bright earth/dark earth/ deep space. We will follow these changes for two orbits with each MAMA with internal lamp and the medium dispersion echelle formats in order to obtain a two-dimension series of reference points on the 2-dimensional detector format. Exposures will be done using the 0.1X0.03 aperture and medium resolution echelle gratings, and will have exposure times of 120 seconds for deep, sharp spectral line images.
For each orbit, six spectral line images will follow each other, then dark frames are interposed with exposure times extending from 300 seconds to 600 seconds. This provides frequent sampling in the portion of the orbit where thermal flexure is largest, while avoiding excessive lamp use when shifts are expected to be slower. The dark frames will also provide a useful addition to the calibration of the MAMA detector dark current.
Note that E140M test is from hot to cold and the E230M test is from cold to hot. If noticeable changes are measured, the complimentary test pair should be considered at a later date.
WFC3 UVIS CCD Daily Monitor
The behavior of the WFC3 UVIS CCD will be monitored daily with a set of full-frame, four-amp bias and dark frames. A smaller set of 2Kx4K subarray biases are acquired at less frequent intervals throughout the cycle to support subarray science observations. The internals from this proposal, along with those from the anneal procedure (11909), will be used to generate the necessary superbias and superdark reference files for the calibration pipeline (CDBS).
Cycle 17: UVIS Bowtie Monitor
Ground testing revealed an intermittent hysteresis type effect in the UVIS detector (both CCDs) at the level of ~1%, lasting hours to days. Initially found via an unexpected bowtie-shaped feature in flatfield ratios, subsequent lab tests on similar e2v devices have since shown that it is also present as simply an overall offset across the entire CCD, i.e., a QE offset without any discernable pattern. These lab tests have further revealed that overexposing the detector to count levels several times full well fills the traps and effectively neutralizes the bowtie. Each visit in this proposal acquires a set of three 3×3 binned internal flatfields: the first unsaturated image will be used to detect any bowtie, the second, highly exposed image will neutralize the bowtie if it is present, and the final image will allow for verification that the bowtie is gone.
UVIS and IR Pointing Stability
This calibration proposal measures the pointing stability of the WFC3 UVIS and IR channels.
Three conditions will be tested: 1) 2-orbit stability after sitting at a constant thermal attitude for 10 orbits 2) 2-orbit stability after sitting at a hot thermal attitude for 10 orbits and then slewing to a cold attitude 3) 2-orbit stability after sitting at a cold thermal attitude for 10 orbits and then slewing to a hot attitude
Stability measurements will be made by a series of short observations of a globular cluster.
FLIGHT OPERATIONS SUMMARY:
Significant Spacecraft Anomalies: (The following are preliminary reports of potential non-nominal performance that will be investigated.)
COMPLETED OPS REQUEST: (None)
COMPLETED OPS NOTES: (None)
FGS GSAcq 05 05
FGS REAcq 12 12
OBAD with Maneuver 02 02
SIGNIFICANT EVENTS: (None)